OpenStudy (anonymous):

Choose the correct simplification of the expression (a^2/b^3)^4 a^6/b^7 b^12 a^8b^15 a^6/b^12

OpenStudy (anonymous):

Notice that\[(\frac{a^2}{b^3})^4=\frac{(a^2)^4}{(b^3)^4}\] Can you simplify from there?

OpenStudy (anonymous):

can you help me please